Two angle are complementary and the measure of the smaller angle is 50° less than the measure of the larger. Find the measure of the larger angle
Answer by ikleyn(53763)

.

x degrees the larger

(x-50) degrees the smaller.


    Their sum must be 90 degrees


        x + (x-50) = 90

          2x       = 90 + 50 = 140  degrees

           x                 = 140/2 = 70 degrees.


ANSWER.  The larger angle is 70 degrees.

Solved.